calculate the value of sin(-45) without using calculator
 

Grade:10

2 Answers

CHALLA NAVNEESWAR REDDY
7 years ago
we know that,
sin(-x) =-sin x
therefore
from sin(-45),
=-sin45      =  we know that  sin 45=1/root 2
therefore  sin(-45)  = -sin45     =-1/root2
Vinay
7 years ago
sin(-45) = sin(0-45)
=> sin(a-b) = sin(a)cos(b) – sin(b)cos(a)
so,
      sin(0-45) = sin(0)cos(45) – sin(45)cos(0)
                    = 0 – (1/sqrt(2))1
                    = -1/sqrt(2)